# Quantum cellular automata and categorical dualities of spin chains

### v2

@@ 1. Introduction
Quantum cellular automata (QCA) are a class of quantum operations that capture the fundamental properties of unitarity and locality [Wat95, Far20, Arr19]. The problem of classifying QCA up to finite depth unitary circuits has been the focus of much attention [GNVW09, SW04, FHH19, FH20, Haa21, Haa22], due to both the fundamental nature of QCA and their wide ranging applications. These include invertible quantum phases of matter [FHH20, JM21], floquet quantum dynamics [GNP21, PFM+16, PFVP17, PVF18, PM17, ZL21, AHLM23], many-body localization [LE24], unitary tensor network operators [CPGSV17, SSBC18, PC20, PTSC21, RWW20], and simulations of quantum field theories [ANW11, FS20, BM20].
@@
We can also use our formalism to make contact with the literature concerning symmetric QCA (sQCA). In this setting, we consider a global, on-site finite group $G$ acting faithfully by unitaries in a self-dual representation. Then the symmetric operators $A^{G}$ have DHR category equivalent to the Drinfeld center $\mathcal{Z}(\text{Rep}(G))$. We can immediately recover an $H^{2}(G,U(1))$ index (c.f. [GSSC20]) as a consequence of our framework (see Theorem 6.3).
@@
In other words, for an onsite group symmetry the DHR action of a duality on the quantum double completely classifies dualities up to finite-depth circuits that are globally symmetric, and have trivial $H^{2}(G,U(1))$ index. In general, the difference between globally symmetric finite-depth circuits with trivial $H^{2}(G,U(1))$ index and locally symmetric finite-depth circuits is unclear. Here, a locally symmetric finite-depth circuit must have some decomposition into local quantum gates that individually commute with the symmetry, whereas a globally symmetric finite-depth circuit only needs to commute with the symmetry globally. Some progress in this direction has been made in [GSSC20], utilizing a generalized GNVW index. Unfortunatey these results do not map precisely onto our framework. However, for abelian groups in the regular representation, it was recently shown that globally symmetric finite depth circuits with trivial $H^{2}(G,U(1))$ index are always locally symmetric [MLC24, Theorem 1]. Corollary 1.4 combined with this result then yields an alternate version of the main result in [MLC24], which uses a different version of the index.
@@
The remainder of this work is presented as follows. In Section 2 we introduce background material about categorical dualities of quantum spin chains. In Section 3 we describe dualities in the setting of fusion spin chains. In Section 4 we describe categorical inclusions defined by module categories over algebra objects. In Section 5 we describe how DHR bimodules can be applied to solve the extension problem for dualities of quantum spin chains. In Section 6 we present examples that generalize the KW duality, and give our applications.
@@ 6. Examples and Applications
For ordinary spin chains, this agrees with the GNVW index [JL24, Proposition 4.2]. Our goal is to show that for the group case, if a $G$-duality $\alpha$ duality admits an extension to a QCA $\widetilde{\alpha}$, then $\text{Ind}(\alpha)=\text{Ind}(\widetilde{\alpha})$. This would imply that if $\text{Ind}(\alpha)=1$, $\widetilde{\alpha}$ is a finite depth circuit.
@@
Proof of Corollary 1.4. Suppose $\alpha:A^{G}\rightarrow A^{G}$ is a G-duality such that $\text{Ind}(\alpha)=1$ and $\text{DHR}(\alpha)=\text{Id}_{\mathcal{Z}(\text{Rep}(G))}$. Then by Theorem 1.3 there is a lift of $\alpha$ to an sQCA $\widetilde{\alpha}$ which can be chosen by 6.3 to have trivial $H^{2}$ index. Furthermore, by Proposition 6.4, $\text{Ind}(\widetilde{\alpha})=1$, so $\widetilde{\alpha}$ is a finite depth circuit.
@@ Acknowledgments
CJ and KS were supported by NSF DMS-2247202. DJW was supported in part by the Australian Research Council Discovery Early Career Research Award (DE220100625). This paper grew out of discussions that took place during the “QCA group” at the workshop “Fusion categories and tensor networks” at the American Institute of Mathematics in 2022. After completing the first arXiv version of this work we became aware of the related work [MLC24]. We have now included comments on the relationship between our results and those of [MLC24].

### v3

@@ 1. Introduction
Quantum cellular automata (QCA) are a class of quantum operations that capture the fundamental properties of unitarity and locality [Wat95, Far20, Arr19]. The problem of classifying QCA up to finite-depth unitary circuits has been the focus of much attention [GNVW09, SW04, FHH19, FH20, Haa21, Haa22], due to both the fundamental nature of QCA and their wide ranging applications. These include invertible quantum phases of matter [FHH20, JM21], floquet quantum dynamics [GNP21, PFM+16, PFVP17, PVF18, PM17, ZL21, AHLM23], many-body localization [LE24a], unitary tensor network operators [CPGSV17, SSBC18, PC20, PTSC21, RWW20], and simulations of quantum field theories [ANW11, FS20, BM20].
@@
We also use our formalism to make contact with the literature concerning symmetric QCA (sQCA). In this setting, we consider a finite gloabl symmetry group $G$ acting faithfully by on-site unitaries in a self-dual representation. Then the symmetric operators $A^{G}$ have DHR category equivalent to the Drinfeld center $\mathcal{Z}(\text{Rep}(G))$. We immediately recover an $H^{2}(G,U(1))$ index (c.f. [GSSC20]) as a consequence of our framework (see Theorem 6.3).
@@
For any finite group $G$, a pair of G-dualities in the regular representation differ by a locally symmetric finite-depth circuit if and only if they have the same numerical index and induce the same braided monoidal autoequivalence (up to natural isomorphism) on $\mathcal{Z}(\text{Rep}(G))$.
@@
The remainder of this work is presented as follows. In Section 2 we introduce background material on categorical dualities of quantum spin chains. In Section 3 we describe dualities in the context of fusion spin chains. In Section 4 we describe categorical inclusions defined by module categories over algebra objects. In Section 5 we describe how DHR bimodules can be applied to solve the extension problem for dualities of quantum spin chains. In Section 6 we present examples that generalize the KW duality, and give our applications.
@@ 6. Examples and Applications
For ordinary spin chains, this agrees with the GNVW index [JL24, Proposition 4.2]. Our goal is to show that for the group case, if a $G$-duality $\alpha$ duality admits an extension to a QCA $\widetilde{\alpha}$, then $\text{Ind}(\alpha)=\text{Ind}(\widetilde{\alpha})$. This would imply that if $\text{Ind}(\alpha)=1$, $\widetilde{\alpha}$ is a finite-depth circuit.
@@
Proof of Corollary 1.4. Suppose $\alpha:A^{G}\rightarrow A^{G}$ is a G-duality such that $\text{Ind}(\alpha)=1$ and $\text{DHR}(\alpha)=\text{Id}_{\mathcal{Z}(\text{Rep}(G))}$. Then by Theorem 1.3 there is a lift of $\alpha$ to an sQCA $\widetilde{\alpha}$ which can be chosen by 6.3 to have trivial $H^{2}$ index. Furthermore, by Proposition 6.4, $\text{Ind}(\widetilde{\alpha})=1$, so $\widetilde{\alpha}$ is a finite-depth circuit.
Now, for abelian $G$ in the regular representation, it was recently shown that globally symmetric circuits with trivial $H^{2}$ invariant are locally symmetric [MLC24, Theorem 1]. More generally, the corresponding statement for groups $G$, which are not necessarily abelian, is shown (implicitly [^1] The argument stated there uses ancilla but can easily be modified to avoid this step by coarse-graining and applying Fell’s absorption principle in the regular representation) in [LE24b]. This concludes the proof.
@@ Acknowledgments
CJ and KS were supported by NSF DMS-2247202. DJW was supported in part by the Australian Research Council Discovery Early Career Research Award (DE220100625). This paper grew out of discussions that took place during the “QCA group” at the workshop “Fusion categories and tensor networks” at the American Institute of Mathematics in 2022. After completing the first arXiv version of this work we became aware of the related work [MLC24]. We have now included comments on the relationship between our results and those of [MLC24]. We also thank Dominic Else for pointing out how the results of [LE24b] can be used to obtain Corollary 1.4, which is a stronger statement than in earlier versions.